Which of the following does NOT represent a mole?
A). 18 g of H2O
B). 40 g of calcium
C). 6.02 x 10^23 atoms of copper
D). 16 grams of oxygen gas (O2)
1
Expert's answer
2018-01-23T15:14:08-0500
For variants A and B the masses are equal Molar mass of the compounds. For variant C 6.02 x 10^23 is the Avogadro’s number, therefore the three variants represent 1 mole.
For variant D:
n(O_2 )=(m(O_2))/(M(O_2))=(16 g)/(32 g/mol)=0.5 mol
So, 16 grams of oxygen do not represent a mole.
Answer:
D